OpenStudy (anonymous):

An average person in the United States throws away 4.5 × 10–1 kg of trash per day. A single gram of trash has an average volume of 3 cm3. How many cubic meters of trash would the average person in the United States produce each year (1 m3 = 1 × 106 cm3)? 4.9 × 10^? m3

OpenStudy (dbzfan836):

@jessicatucker your calculations were quite off, yet i will tell you step by step how i did the problem.

OpenStudy (dbzfan836):

first, I took your original problem, 4.5 * 10 - 1, and calculated it like so: 4.5*10=45 45-1=44 now the cubic meters, are this: 44*1000=44,000 44,000*3=132,000 Now to make it into meters: 132,000 / 100

OpenStudy (dbzfan836):

and it equals: you solve it

OpenStudy (dbzfan836):

Earlier i didnt give you the right answer, due to me making a mistake.

OpenStudy (ibbutibbu.):

@Dbzfan836 lawlz